# Statistical features of quantum chaos using the Krylov operator complexity

### v1

@@ Appendix A Check the normal distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=100$
Here we show the data fitting of the histogram in Fig. 4 to the normal distribution. For GOE and GUE, the fitted distribution is shown in Fig. 6. For uniform distributions, the fitted distribution is shown in Fig. 7. The average value and the standard deviation are listed in Table 1, where we compare the values obtained from the sampled data and from the fitted normal distribution.
@@
$\mathcal{N}_{max}=100$
$\mu_{0}$ $\sigma_{0}$
nonchaos GOE Fit 0.08205 0.00125033
Data 0.0821079 0.00124627
GUE Fit 0.0818423 0.000799776
Data 0.0818522 0.000818971
URE Fit 0.0819847 0.00122253
Data 0.0819877 0.00122892
UIM Fit 0.0822362 0.001221
Data 0.0822584 0.0012246
UCP Fit 0.0818089 0.000796413
Data 0.0817893 0.000813977
chaos GOE Fit 0.000360084 0.0000311866
Data 0.000361536 0.0000312431
GUE Fit 0.000239301 0.0000196699
Data 0.00024081 0.0000194523
URE Fit 0.000313361 0.0000277979
Data 0.000315455 0.0000278845
UIM Fit 0.000314084 0.0000270829
Data 0.000315694 0.0000267367
UCP Fit 0.000210981 0.0000167099
Data 0.000211866 0.0000168488
Appendix B Different choice of algorithm step sizes for $b_{n}$
For the overlapping-separating behavior, here we show its independence on the sets of coefficients used in computing $\sigma^{2}$. Now we choose the Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$.
Firstly, we compute the variance $\sigma^{2}$ (2.7) for GOE and GUE, where the first $1000$ samples are shown in Fig. 8. We see that GOE and GUE mix together in the nonchaotic case and separate from each other in the chaotic case.
(a) Sampled variances $\sigma^{2}$ of GOE and GUE in chaotic case $a=1$.
(b) Sampled variances $\sigma^{2}$ of GOE and GUE in nonchaotic case $a=0$.
Figure 8: The first $1000$ samples of variances $\sigma^{2}$ for GOE and GUE. Here the Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$ are used.
Then, the histogram of all $5000$ samples are shown in Fig. 9. Obviously they still resemble normal distributions and are overlapped in the nonchaotic case and well-separated in the chaotic case.
(a) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
(b) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
Figure 9: The histogram of variances $\sigma^{2}$ after sampling the initial operator 5000 times from GOE and GUE. The Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$ are used. They are overlapped in the nonchaotic case and well-separated in the chaotic case.
Finally, the fitted normal distributions are shown in Fig. 10. So we still get the resulting normal distributions and the overlapping-splitting behaviors, although the detailed values of the average and the standard deviation change.
(c) Chaotic case $a=1$ of GOE. Skewness=0.107514, Kurtosis=3.01354.
(d) Chaotic case $a=1$ of GUE. Skewness=0.0881284, Kurtosis=3.06162.
x
Appendix C Check the normal distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=5$
Here we study the case of $\mathcal{N}_{max}=5$. Then the sets of coefficients $b_{n}$ are chosen from step-sizes $151<n<$ 15. We repeat the computation for GOE and GUE, and show that there is neither normal distributions nor separating behaviors.
The histogram of all $5000$ samples are shown in Fig. 11. Obviously they are not normal distributions and are not separated in the chaotic case.
(a) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
(b) Histogram of $\sigma^{2}$ of GOE and GUE in the nonchaotic case $a=0$.
Figure 11: The histogram of variances $\sigma^{2}$ after sampling the initial operator 5000 times from GOE and GUE, $\mathcal{N}_{max}=5$. Obviously they are not normal distributions.
Anyway, if we force to fit the histogram to normal distributions, the results are shown Fig. 12. From the ridiculous value of skewness and kurtosis, we understand that the histogram can not be the normal distribution. Also, from the values listed in Table 2, we see that there is large discrepancy between values from the sample data and values from the fitted distribution. So for $\mathcal{N}_{max}=5$, the resulting distributions are not normal and there is no separating behaviors between GOE and GUE.
(a) $N=5$ chaos, Gaussian U ensembles. Skewness=1.19843, Kurtosis=5.18562.
(b) $N=5$ nonchaos, Gaussian U ensembles. Skewness=4.30801, Kurtosis=50.7359.
(c) $N=5$ chaos, Gaussian O ensembles. Skewness=1.6231, Kurtosis=7.05429.
(d) $N=5$ nonchaos, Gaussian O ensembles. Skewness=5.83862, Kurtosis=73.6201.
Figure 12: Data fitting for GOE and GUE, $\mathcal{N}_{max}=5$. The large value of skewness and kurtosis means that they can not be normal.
$\mathcal{N}_{max}=5$
$\mu_{0}$ $\sigma_{0}$
nonchaos GOE Fit 0.128068 0.099812
Data 0.188669 0.178039
GUE Fit 0.0967919 0.0465464
Data 0.119184 0.0752851
URE Fit 0.119756 0.0814372
Data 0.169944 0.146747
UIM Fit 0.206881 0.220034
Data 0.467013 0.602757
UCP Fit 0.0678911 0.0397609
Data 0.0877107 0.0616032
chaos GOE Fit 0.127619 0.0872543
Data 0.170123 0.112041
GUE Fit 0.104014 0.0521167
Data 0.121707 0.0610372
URE Fit 0.112995 0.0772997
Data 0.150244 0.0991289
UIM Fit 0.20724 0.0973048
Data 0.253492 0.14551
UCP Fit 0.0688293 0.0407356
Data 0.0853616 0.0509862
Table 2: Average $\mu_{0}$ and standard deviation $\sigma_{0}$ for the data fitting, $\mathcal{N}_{max}=5$. There is large discrepancy between values computed from the sampled data and values of the fitted normal distribution.

### v2

@@ Appendix A Check the normal distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=100$
Here we show the data fitting of the histogram in Figure 7 to the normal distribution. For GOE and GUE, the fitted distribution is shown in Figure 10. For uniform distributions, the fitted distribution is shown in Figure 11. The average value and the standard deviation are listed in Table 1, where we compare the values obtained from the sampled data and from the fitted normal distribution.
Appendix B Different choice of algorithm step sizes for $b_{n}$
For the overlapping-separating behavior, here we show its independence on the sets of coefficients used in computing $\sigma^{2}$. Now we choose the Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $10\mathcal{N}_{max}\leq n\leq 15\mathcal{N}_{max}$ of Figure 4.
Firstly, we compute the variance $\sigma^{2}$ (3.3), where the first $1000$ samples are shown in Figure 12. We see that they mix together in the nonchaotic case and separate from each other in the chaotic case.
Then, the histogram of all $5000$ samples are shown in Figure 13. Obviously they still resemble normal distributions and are overlapped in the nonchaotic case and well-separated in the chaotic case.
Finally, the fitted normal distributions are shown in Figure 14 and 15. So we still get the resulting normal distributions and the overlapping-splitting behaviors, although the detailed values of the average and the standard deviation change.
Appendix C Check the chi-square distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=5$
Here we study the case of $\mathcal{N}_{max}=5$. Then the sets of coefficients $b_{n}$ are chosen from step-sizes $151<n<$ 15. We repeat the computation and show that there is neither normal distributions nor separating behaviors.
The histogram of all $5000$ samples are shown in Figure 16. Obviously they are not normal distributions and are not separated in the chaotic case. So the normal distribution is necessary for the overlapping-separating behavior.
We fit the histogram to the rescaled chi-square distribution, the results are shown in Figure 17 for the Gaussian ensembles and in Figure 18 for the uniform distributions,. We see that the rescaled chi-square distribution indeed captures the resulting distribution $f_{\sigma^{2}}$, so it verifies our analysis in Section 3.3.
@@ Appendix C Check the chi-square distribution of $\sigma^{2}$ for $\mathcal{N}_{max}=5$
(a) Chaotic case $a=1$ of UIM. Skewness=0.225317, Kurtosis=3.18781.
(b) Chaotic case $a=1$ of URE. Skewness=0.131034, Kurtosis=3.00302.
(c) Chaotic case $a=1$ of UCP. Skewness=0.167094, Kurtosis=3.06099.
(d) Nonchaotic case $a=0$ of UIM. Skewness=0.1411, Kurtosis=3.08718.
(e) Nonchaotic case $a=0$ of URE. Skewness=0.0823458, Kurtosis=2.97123.
(f) Nonchaotic case $a=0$ of UCP. Skewness=-0.0801398, Kurtosis=3.04548.
$\mathcal{N}_{max}=100$
$\mu_{0}$ $\sigma_{0}$
nonchaos GOE Fit 0.08205 0.00125033
Data 0.0821079 0.00124627
GUE Fit 0.0818423 0.000799776
Data 0.0818522 0.000818971
URE Fit 0.0819622 0.0012411
Data 0.0819846 0.0012243
UIM Fit 0.0821883 0.00122778
Data 0.0822335 0.00124233
UCP Fit 0.0817821 0.000810266
Data 0.0817644 0.000816963
chaos GOE Fit 0.000360084 0.0000311866
Data 0.000361536 0.0000312431
GUE Fit 0.000239301 0.0000196699
Data 0.00024081 0.0000194523
URE Fit 0.000314526 0.0000268396
Data 0.000315565 0.0000270371
UIM Fit 0.000314332 0.0000278555
Data 0.000315618 0.0000277744
UCP Fit 2.11016 0.163631
Data 2.11775 0.164331
(a) Sampled variances $\sigma^{2}$ of GOE and GUE in chaotic case $a=1$.
(b) Sampled variances $\sigma^{2}$ of GOE and GUE in nonchaotic case $a=0$.
Figure 12: The first $1000$ samples of variances $\sigma^{2}$ for GOE, GUE, URE, UIM and UCP. Here the Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$ are used.
(a) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
(b) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
Figure 13: The histogram of variances $\sigma^{2}$ after sampling the initial operator 5000 times from GOE, GUE, URE, UIM and UCP. The Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$ are used. They are overlapped in the nonchaotic case and well-separated in the chaotic case.
(c) Chaotic case $a=1$ of GOE. Skewness=0.107514, Kurtosis=3.01354.
(d) Chaotic case $a=1$ of GUE. Skewness=0.0881284, Kurtosis=3.06162.
(a) Chaotic case $a=1$ of UIM. Skewness=0.160477, Kurtosis=3.02122.
(b) Chaotic case $a=1$ of URE. Skewness=0.200908, Kurtosis=3.02255.
(c) Chaotic case $a=1$ of UCP. Skewness=0.181879, Kurtosis=3.18625.
(d) Nonchaotic case $a=0$ of UIM. Skewness=-0.165951, Kurtosis=3.02278.
(e) Nonchaotic case $a=0$ of URE. Skewness=-0.0222373, Kurtosis=2.87503.
(f) Nonchaotic case $a=0$ of UCP. Skewness=-0.169296, Kurtosis=3.0196.
Figure 15: Fitted normal distribution for uniform distributions, $\mathcal{N}_{max}=100$. The Lanczos coefficients $\left\{b_{n}\right\}$ from step-sizes $1000\leq n\leq 1500$ are used.
(a) Histogram of $\sigma^{2}$ of GOE and GUE in the chaotic case $a=1$.
(b) Histogram of $\sigma^{2}$ of GOE and GUE in the nonchaotic case $a=0$.
Figure 16: The histogram of variances $\sigma^{2}$ after sampling the initial operator 5000 times from GOE, GUE, URE, UIM and UCP, with $\mathcal{N}_{max}=5$. Obviously they are not normal distributions.
(a) $N=5$ chaos, GUE. Skewness=1.19843, Kurtosis=5.18562.
(b) $N=5$ nonchaos, GUE. Skewness=4.30801, Kurtosis=50.7359.
(c) $N=5$ chaos, GOE. Skewness=1.6231, Kurtosis=7.05429.
(d) $N=5$ nonchaos, GOE. Skewness=5.83862, Kurtosis=73.6201.
Figure 17: Data fitting of rescaled chi-square distribution for GOE and GUE, $\mathcal{N}_{max}=5$.
@@
Figure 18: Data fitting of rescaled chi-square distribution for uniform distributions, $\mathcal{N}_{max}=5$.